OpenStudy (anonymous):

Let f(x)=x^3-2x+1,g(x)=-2x and x=1 Find the area between the bounded regions

OpenStudy (anonymous):

How can u attach a picture ?

OpenStudy (anonymous):

first find the point of intersection (set the two functions equal solve for x) x^3-2x+1=-2x x^3+1=0 (x+1)(x^2-x+1)=0 x+1=0 (x^2-x+1)=0 x=-1 (neglect the two imaginary roots given by (x^2-x+1)=0) area is given by \[\Large \int\limits_{-1}^{1}(x^3-2x+1+2x)dx\]

OpenStudy (anonymous):

Then you do integration for the numbers and that's the answer ?

OpenStudy (anonymous):

yes .

OpenStudy (anonymous):

Ok let me work it out and ill get back to you Thanks

OpenStudy (anonymous):

yw !

OpenStudy (anonymous):

Is the smallest number always a on the integral?

OpenStudy (anonymous):

Which line is the one on top? X^3-2x+1 or -2x

OpenStudy (anonymous):

the x^3-2x+1 is on the top.

OpenStudy (anonymous):

O got 1/2 for my answer is it correct ?

OpenStudy (anonymous):

I *